# Spontaneous Magnon Decays from Nonrelativistic Time-Reversal Symmetry Breaking in Altermagnets

### v1

@@
Spontaneous Magnon Decays from Nonrelativistic Time-Reversal Symmetry Breaking in Altermagnets
Rintaro Eto Department of Applied Physics, Waseda University, Okubo, Shinjuku-ku, Tokyo 169-8555, Japan Institut für Physik, Johannes Gutenberg Universität Mainz, D-55099 Mainz, Germany    Matthias Gohlke Theory of Quantum Matter Unit, Okinawa Institute of Science and Technology Graduate University, Onna-son, Okinawa 904-0495, Japan Max Planck Institute for the Physics of Complex Systems, Nöthnitzer Str. 38, 01187 Dresden, Germany    Jairo Sinova Institut für Physik, Johannes Gutenberg Universität Mainz, D-55099 Mainz, Germany    Masahito Mochizuki Department of Applied Physics, Waseda University, Okubo, Shinjuku-ku, Tokyo 169-8555, Japan    Alexander L. Chernyshev Department of Physics and Astronomy, University of California, Irvine, California 92697, USA    Alexander Mook Institut für Physik, Johannes Gutenberg Universität Mainz, D-55099 Mainz, Germany
(February 27, 2025)
@@
The concept of a quasiparticle—an emergent, long-lived excitation that weakly interacts with its environment—is foundational to condensed matter physics, shaping our understanding of diverse systems [1, 2]. A central question is when and how quasiparticles break down at zero temperature due to quantum fluctuations [3, 4]. In bosonic systems, many-body interactions can trigger spontaneous quasiparticle decay, as seen in the instability of Bogoliubov excitations [5, 6] and anharmonic phonon decay in solids [7]. In quantum magnets, magnons—collective spin-wave excitations—are subject to unavoidable nonlinear interactions, leading to pronounced decay in frustrated and low-dimensional systems [3]. Understanding these decay mechanisms is crucial for determining whether quasiparticles remain the appropriate low-energy degrees of freedom or give way to more complex excitations. Experimentally, such decay manifests as spectral broadening and weight redistribution in neutron scattering and optical spectroscopy, offering insight into a material’s quantum many-body physics.
@@
In many magnetically ordered insulators, spin space remains approximately isotropic and their leading-order physics is described by the paradigmatic Heisenberg model. This is the case when relativistic effects are negligible, such that spin space and real space are not connected. Once the spontaneous magnetic long-range order breaks continuous spin-rotation symmetry, gapless Nambu-Goldstone modes and universal low-energy behavior emerge. In collinear magnets—the simplest form of magnetic order—a dichotomy of magnon stability emerges: ferromagnetic magnons are inherently stable, as their ground state and excitations are exact eigenstates free of quantum fluctuations, whereas antiferromagnetic magnons experience fluctuations that can induce spontaneous decay [8, 3]. This dichotomy collapses when spin space becomes connected to real space by the relativistic spin-orbit coupling, noncollinear order, or dipolar interactions, making magnon decay a more intricate problem [9, 10, 11, 3, 12, 13, 14, 15, 16, 17, 18].
@@
Altermagnets—a newly identified class of collinear magnets [19, 20]—challenge this dichotomy even within the nonrelativistic limit. Like antiferromagnets, they feature compensated magnetic order and quantum fluctuations, yet they also break time-reversal symmetry and exhibit nonrelativistic band splitting reminiscent of ferromagnets. Their unconventional beyond-s-wave even-parity spin splitting (d-, g-, or i-wave) of magnons [21, 22] raises a fundamental question: Can altermagnets support stable magnons, and how does their distinctive spin splitting interact with quantum many-body fluctuations to influence magnon stability? This question has gained urgency following quantum simulations of spin-$1/2$ altermagnets [23], which suggest pronounced nonclassical effects and high-energy magnon instability.
@@
Figure 1: Stability of low-energy magnons in the three classes of collinear magnets in the nonrelativistic limit of an isotropic Heisenberg model. Due to spin conservation, magnons carry a spin quantum number ±$\pm 1$, indicated by color, and potential decay processes have to conserve spin. (a) Class-(i) magnets: Magnets with s-wave spin splitting, such as ferromagnets and ferrimagnets, exhibit inherently stable magnons at low energies. In ferromagnets, there are no quantum fluctuations and there is only one (acoustic) magnon branch, which can be considered the limit of infinite s-wave splitting. In contrast, ferrimagnets fluctuate. They support two oppositely polarized sublattices with unequal moments, and their spin splitting, i.e., the gap $\Delta_{0}$, is finite. It prohibits magnon decay at energies below $\Delta_{0}$, ensuring their stability. The indicated decay process of the magenta magnon is only possible for high-energy magnons above $\Delta_{0}$, such that one decay product can carry opposite spin (cyan). Thus, the lower magnon branch is stable below the gap $\Delta_{0}$, but the upper magnon branch can decay. (b) Class-(ii) magnets: Antiferromagnets fluctuate and have spin-degenerate magnons. Stability depends on the curvature of the magnon spectrum: negatively curved spectra ($\kappa<0$) lead to stable magnons, while positively curved spectra ($\kappa>0$) result in instability. (c) Class-(iii) magnets: Altermagnets are characterized by unconventional time-reversal symmetry breaking and beyond-s-wave even-parity spin splitting such as d/g/i-wave. In nodal planes, i.e., directions without spin splitting, the stability mirrors that of class-(ii) magnets, where the curvature of the spectrum dictates whether magnons can decay, but is additionally modified by the altermagnetic splitting, as indicated by the asterisk. Along directions with finite spin splitting the upper magnon branch is universally unstable due to the indicated decay process. The lower magnon branch exhibits rich stability physics that depends on the symmetry of the splitting as indicated by the asterisk: In d-wave magnets, it is either completely stable for $\kappa<0$ or exhibits a d-wave-shaped island of stability in the vicinity of the origin for $\kappa>0$. In contrast, there is no such island of stability for general g-wave and i-wave altermagnets.
@@
For insulators with localized electronic degrees of freedom, all three classes of magnets adhere to the nonrelativistic SO$(3)$-symmetric Heisenberg model ($\hbar=1$),
@@
where ${\boldsymbol{S}}_{\boldsymbol{r}}$ is a local spin operator of length $S$ at site ${\boldsymbol{r}}$, and $J_{{\boldsymbol{r}},{\boldsymbol{r}}^{\prime}}$ the exchange constant between spins at ${\boldsymbol{r}}$ and ${\boldsymbol{r}}^{\prime}$. In the collinear ground state, spontaneous symmetry breaking reduces SO$(3)$ to SO$(2)$, with the ordered moment chosen along the $z$-axis. As a result, Goldstone modes emerge, and the total spin component $S^{z}=\sum_{\boldsymbol{r}}S_{\boldsymbol{r}}^{z}$ remains a good quantum number. Single spin-flip excitations (magnons) thus carry spin ±$\pm 1$ (also denoted by $\uparrow$ and $\downarrow$) relative to the ground state, and follow dispersions $\epsilon_{\pm}({\boldsymbol{k}})$, where ${\boldsymbol{k}}$ is the crystal momentum, and “$\pm$” indicates spin, not necessarily energetic order. Optical magnon branches are neglected as they do not impact low-energy physics.
@@
Class-(i) magnets: Ferromagnets inherently support stable magnons because their ground state is an exact eigenstate. Ferrimagnets, however, fluctuate because their sublattice-polarized ground state is not an eigenstate. Having two antiparallel but uncompensated sublattice magnetizations, they feature two magnon branches: a gapless branch, $\epsilon_{-}({\boldsymbol{k}})=A|{\boldsymbol{k}}|^{2}$, and a gapped branch, $\epsilon_{+}({\boldsymbol{k}})=A|{\boldsymbol{k}}|^{2}+\Delta_{0}$, where $A>0$ is the spin-wave stiffness and $\Delta_{0}>0$ the gap, see Fig. 1a. Here, s-wave spin splitting manifests as an isotropic energy difference: $\epsilon_{+}({\boldsymbol{k}})-\epsilon_{-}({\boldsymbol{k}})=\Delta_{0}$. The positive curvature of the gapless branch implies that any initial magnon could theoretically decay into three magnons at smaller momenta while conserving energy and momentum. However, Eq. (2) requires at least one decay product to have opposite spin, precluding decay into the gapless branch alone. Decay becomes possible only at energies exceeding $\Delta_{0}$, where one decay product can transition into the gapped branch. Consequently, the gapless magnon branch in class-(i) magnets is universally stable below the energy gap $\Delta_{0}$ and protected by s-wave spin splitting. The gapped branch, however, can always decay.
@@
Class-(iii) magnets: To derive general insights into magnon decay in altermagnets, we perform a low-energy expansion of the magnon dispersion in two dimensions. For simplicity, we assume the dispersion to be spatially isotropic, allowing crystallographic symmetries to enter only through the altermagnetic splitting. Under these approximations, the magnon dispersion takes the form
@@
where $v>0$ is the magnon velocity, $\kappa$ represents the leading (cubic) non-altermagnetic nonlinearity, $\tan\varphi_{{\boldsymbol{k}}}=k_{y}/k_{x}$, and $\Delta\geq 0$ parametrizes the altermagnetic splitting with a d-wave ($n=2$), g-wave ($n=4$) or i-wave character ($n=6$). The unconventional time-reversal symmetry breaking results in an even-parity spin splitting gap $\epsilon_{+}({\boldsymbol{k}})-\epsilon_{-}({\boldsymbol{k}})=2\Delta\sin\left(n\varphi_{{\boldsymbol{k}}}\right)|{\boldsymbol{k}}|^{n}$ that modulates with $n\varphi_{{\boldsymbol{k}}}$. Along a nodal plane (or line in 2D), e.g., ${\boldsymbol{k}}=(k_{x},k_{y})=(k,0)$, the splitting vanishes, and the degenerate spectrum suggests that the curvature argument relevant to class-(ii) magnets will also play a role here. Away from the nodal planes, the magnon bands split, opening up novel decay channels, see Fig. 1c.
We begin with discussing magnon stability along the nodal planes. If $\kappa>0$, nodal-plane magnons can decay similarly to magnons in class-(ii) magnets. This leaves $\kappa<0$ as the relevant case, where magnons in class-(ii) magnets remain stable. Consider the case depicted in Fig. 2a, where a nodal plane magnon decays into a magnon close to the Goldstone mode, and two lower-branch magnons to the left and right of the nodal plane. As shown in the Supplementary Note 1, this process obeys Eq. (2), if $\Delta>\Delta^{\star}$, where
$\Delta^{\star}=\sqrt{3}\frac{2^{n-5/2}}{nk^{n-2}}\sqrt{-\kappa}\sqrt{4v+3k^{2}\kappa}\quad\text{for}\quad\kappa<0.$ (4)
In the long-wavelength limit ($k\to 0$), $\Delta^{\star}$ diverges for $n=4$ (g-wave) and $n=6$ (i-wave) due to the $k^{n-2}$ factor in the denominator. This indicates that nodal-plane magnons are stable in the vicinity of the origin. For $n=2$ (d-wave), however, $\Delta^{\star}$ converges to
@@
Thus, d-wave magnets with a negatively curved magnon spectrum ($\kappa<0$) may still exhibit unstable nodal-plane magnons if the altermagnetic splitting exceeds $\Delta^{\star}_{\text{d-wave}}$.
Next, we analyze the decay of the lower-branch magnon along a general direction, where it could decay into two lower-branch magnons and a magnon with opposite spin close to the Goldstone mode, as shown in Fig. 2b. For $\kappa<0$, this decay process is generally kinematically forbidden. For $\kappa>0$, as shown in Supplementary Note 1, it remains forbidden if
@@
Thus, in d-wave altermagnets ($n=2$), lower-branch magnons at $|{\boldsymbol{k}}|<k^{\star}_{\text{d-wave}}\left|\sin\left(2\varphi_{{\boldsymbol{k}}}\right)\right|$, with
$k^{\star}_{\text{d-wave}}=\frac{2\Delta}{3\kappa}\quad\text{for}\quad\kappa>0$ (7)
are stable. No such long-wavelength solution can be found for g-wave magnets ($n=4$) and i-wave magnets ($n=6$) because of the negative power of $k^{3-n}$, which turns the inequality into a statement about short wavelengths, where the approximation in Eq. (3) is no longer valid. Thus, the altermagnetic splitting of g-wave and i-wave magnets does not protect the lower branch from decays and it is the dominating cubic nonlinearity that decides the fate of their lower-branch magnons.
These qualitative kinematic considerations for the lower branch can be summarized as follows. Two cases must be distinguished for d-wave magnets: For $\kappa<0$ and $\Delta<\Delta^{\star}_{\text{d-wave}}$, the lower magnon branch remains stable. For $\kappa<0$ and $\Delta>\Delta^{\star}_{\text{d-wave}}$ or for $\kappa>0$, it exhibits a d-wave-shaped “island of stability” around the origin, with nodes at the nodal planes. The extent of this stability region is given by $k^{\star}_{\text{d-wave}}$ in Eq. (7). In contrast, in g-wave and i-wave altermagnets, its stability depends solely on the sign of $\kappa$, as in class-(ii) magnets. The qualitative distinction between d-wave and g,i-wave altermagnets arises because the altermagnetic splitting in d-wave magnets depends on a lower power of momentum than the leading (cubic) nonaltermagnetic nonlinearity, $\kappa$. In contrast, for g-wave and i-wave altermagnets, the altermagnetic splitting becomes irrelevant in the long-wavelength limit compared to the cubic nonlinearities  [^1] In fine-tuned systems with $\kappa=0$, which we do not explore here, quintic terms, $\sim|{\boldsymbol{k}}|^{5}$, become the dominant nonaltermagnetic nonlinearity, and the quartic g-wave altermagnetic splitting becomes relevant in the long-wavelength limit. Similarly, if both cubic and quintic terms vanish, septic nonlinearities, $\sim|{\boldsymbol{k}}|^{7}$, get dominated by the sextic i-wave altermagnetic splitting..
We now turn to the upper magnon branch, for which two decay channels are indicated in Fig. 2c and Fig. 2d. In the process in Fig. 2c, an upper-branch magnon decays into a mode close to the origin and two modes to the left and right. The process is possible in d-wave magnets even for $\kappa<0$ because their altermagnetic splitting dominates over the cubic nonlinearity at long wavelengths. In contrast, g-wave and i-wave altermagnets do not support this process. However, the process in Fig. 2d has the upper-branch magnon decay into a lower-branch magnon at approximately the same momentum and two upper-branch magnons around the origin. As this decay is always kinematically possible, the upper magnon branch in altermagnets is universally unstable.
@@
Figure 3: Nonlinear spin-wave analysis of a two-dimensional d-wave altermagnet. (a) Schematic of the spin model for a two-dimensional d-wave altermagnet on a checkerboard lattice, described by the Hamiltonian in Eq. (8). The system hosts a collinear $(\pi,\pi)$ Néel order with nearest-neighbor (white bonds) and next-nearest-neighbor (red and blue bonds) exchange interactions $J_{1}$ and $J_{2}\pm\Delta$, respectively. The checkerboard modulation—parametrized by $\Delta$—arises from nonmagnetic atoms (silver spheres) that alter the diagonal exchange, inducing a $d$-wave spin splitting of magnons. (b) Diagrammatic representation of the three-magnon scattering process, the leading mechanism for magnon decay. Due to momentum conservation (up to a reciprocal lattice vector), the sum of the momenta of the three intermediate states equals the initial momentum ${\boldsymbol{k}}$. Total spin conservation ensures that an initial single spin-$\uparrow$ magnon scatters into two spin-$\uparrow$ magnons and one spin-$\downarrow$ magnon, with a similar process for an initial spin-$\downarrow$ magnon. (c-f) Renormalized magnon spectrum (lines with dot markers) compared to the linear magnon spectrum (solid lines) for $J_{1}=1$, $J_{2}=-0.5$, $S=1/2$, and selected values of $\Delta$ as indicated. The altermagnetic splitting is most pronounced at $(\frac{\pi}{2},\pm\frac{\pi}{2})$ and zero along the nodal planes, e.g., in the direction from $(0,0)$ to $(\pi,0)$. (g-j) Magnon damping $\Gamma_{\pm}({{\boldsymbol{k}}})$, highlighting the instability of the upper magnon branch. The inset in panel (g) shows the magnetic Brillouin zone with indicated high-symmetry paths. The color of these paths in used on the horizontal momentum axes in panels (c-j) to indicate direction. The cyan/magenta checkerboard pattern indicates the sign of the d-wave altermagnetic spin splitting. The legend in panel (j) distinguishes between nonlinear spin-wave (NLSW) and linear spin-wave (LSW) calculations.
@@
To expand upon the preceding qualitative discussion of magnon decay we consider a d-wave altermagnet on the checkerboard lattice (see Fig. 3a), for which we carry out a nonlinear spin-wave analysis. The spin Hamiltonian under investigation is given by
@@
Here, $J_{1}>0$ is the antiferromagnetic nearest-neighbor, and $J_{2}$ the second-nearest-neighbor exchange interaction. $\Delta>0$ parametrizes the d-wave altermagnetic magnon band splitting. Its sign alternates between white and black squares of the checkerboard, as indicated in Fig. 3a.
Using a $1/S$ spin-wave expansion around the Néel-ordered ground state based on the Holstein-Primakoff transformation [25], we derive the noninteracting single-magnon spectrum $\epsilon_{\pm}({\boldsymbol{k}})=S\sqrt{A({\boldsymbol{k}})^{2}-B({\boldsymbol{k}})^{2}}\pm S\Delta({\boldsymbol{k}})$, with $A({\boldsymbol{k}})$, $B({\boldsymbol{k}})$, and $\Delta({\boldsymbol{k}})$ given in Methods. At low energies, the spectrum can be mapped onto Eq. (3) up to a rotation of the nodal planes and an additional directional dependence of the cubic nonlinearits. Up to order $1/S^{2}$, the renormalized single-magnon spectrum is given by $\tilde{\epsilon}_{\pm}({\boldsymbol{k}})=\epsilon_{\pm}({\boldsymbol{k}})+\delta\epsilon_{\pm}({\boldsymbol{k}})-\mathrm{i}\Gamma_{\pm}({\boldsymbol{k}})$. The real $\delta\epsilon_{\pm}({\boldsymbol{k}})$ corrects the magnon energy and the magnon lifetime $\tau_{\pm}({\boldsymbol{k}})=-1/\mathrm{Im[\tilde{\epsilon}_{\pm}({\boldsymbol{k}})]}$ derives from the decay rate
@@
×$\times\delta\left(\epsilon_{\pm}({\boldsymbol{k}})-\epsilon_{\pm}({\boldsymbol{p}})-\epsilon_{\pm}({\boldsymbol{q}})-\epsilon_{\mp}({\boldsymbol{k}}-{\boldsymbol{p}}-{\boldsymbol{q}})\right),$ (9)
@@
associated with the imaginary part of the self-energy diagram in Fig. 3b, where one encounters the energy conservation condition of Eq. (2). In Eq. (9), $N_{\textrm{muc}}$ is the number of magnetic unit cells, the momenta sums run over the magnetic Brillouin zone, and $W_{\pm}({\boldsymbol{k}},{\boldsymbol{p}},{\boldsymbol{q}})$ is the one-in-three-out interaction vertex (see Methods and Supplementary Note 2).
@@
The interacting magnon spectrum $\tilde{\epsilon}_{\pm}({\boldsymbol{k}})$ is shown in Figs. 3c-j for parameters realizing $\kappa<0$. It exhibits clear d-wave splitting, which increases as $\Delta$ increases from Fig. 3c to Fig. 3f. The many-body interactions cause a stiffening of both bands and, importantly, a damping of the upper magnon branch, see Figs. 3g-j. The damping grows with increasing $\Delta$ and is most pronounced in the short-wavelength limit. Since $\Delta>\Delta^{\star}_{\text{d-wave}}$ in Figs. 3h, the magnon damping is also finite along the nodal planes, that is, along the direction from $(0,0)$ to $(\pi,0)$, in agreement with the kinematic predictions.
The main features of $\Gamma_{\pm}({\boldsymbol{k}})$ in Figs. 3g-j can be understood by setting the interaction vertex in Eq. (9) to one, suggesting that $\Gamma_{\pm}({\boldsymbol{k}})\propto\mathcal{D}_{\pm}(\epsilon_{\pm}({\boldsymbol{k}}),{\boldsymbol{k}})$, where

### v2

@@
Abstract
@@ Introduction
The concept of a quasiparticle—an emergent, long-lived excitation that weakly interacts with its environment—is foundational to condensed matter physics, shaping our understanding of diverse systems [1, 2]. A central question is when and how quasiparticles break down at zero temperature due to quantum fluctuations [3, 4]. In bosonic systems, many-body interactions can trigger spontaneous quasiparticle decay, as seen in the instability of Bogoliubov excitations [5, 6] and anharmonic phonon decay in solids [7]. In quantum magnets, magnons—collective spin-wave excitations—experience nonlinear interactions, leading to decay in frustrated and low-dimensional systems [3]. Understanding these decay mechanisms is crucial for determining whether quasiparticles remain the appropriate low-energy degrees of freedom or give way to more complex excitations. Experimentally, such decay manifests as spectral broadening and weight redistribution in neutron scattering and optical spectroscopy, offering insight into a material’s quantum many-body physics.
@@
In many magnetically ordered insulators, spin space remains approximately isotropic and their leading-order physics is described by the paradigmatic Heisenberg model. This is the case when relativistic effects are negligible, such that spin space and real space are not connected. Spontaneous magnetic long-range order breaks continuous spin-rotation symmetry, leading to gapless Nambu-Goldstone modes and universal low-energy behavior. In collinear magnets—the simplest form of magnetic order—a dichotomy of magnon stability emerges: ferromagnetic magnons are inherently stable, as their ground state and excitations are exact eigenstates free of quantum fluctuations, whereas antiferromagnetic magnons experience fluctuations that can induce spontaneous decay [8, 3]. This dichotomy collapses when spin space becomes connected to real space by the relativistic spin-orbit coupling, noncollinear order, or dipolar interactions, making magnon decay more intricate [9, 10, 11, 3, 12, 13, 14, 15, 16, 17, 18].
@@
Altermagnets—a newly identified class of collinear magnets [19, 20]—challenge this dichotomy even within the nonrelativistic limit. Like antiferromagnets, they feature compensated magnetic order and quantum fluctuations, yet they also break time-reversal symmetry and exhibit nonrelativistic band splitting reminiscent of ferromagnets. Their unconventional beyond-s-wave even-parity spin splitting (d-, g-, or i-wave) of magnons [21, 22, 23, 24] raises a fundamental question: Can altermagnets support stable magnons, and how does their distinctive spin splitting interact with quantum many-body fluctuations to influence magnon stability? This question has gained urgency following quantum simulations of spin-$1/2$ altermagnets [25, 26], which suggest pronounced nonclassical effects and high-energy magnon instability.
@@
Figure 1: Stability of low-energy magnons in the three classes of collinear magnets in the nonrelativistic limit of an isotropic Heisenberg model. Due to spin conservation, magnons carry a spin quantum number $\pm 1$, indicated by color, and potential decay processes have to conserve spin. (a) Class-(i) magnets: Magnets with s-wave spin splitting, such as ferromagnets and ferrimagnets, exhibit inherently stable magnons at low energies. In ferromagnets, there are no quantum fluctuations and there is only one (acoustic) magnon branch, which can be considered the limit of infinite s-wave splitting. In contrast, ferrimagnets fluctuate. They support two oppositely polarized sublattices with unequal moments, and their spin splitting, i.e., the gap $\Delta_{0}$, is finite. It prohibits magnon decay at energies below $\Delta_{0}$, ensuring their stability. The indicated decay process of the magenta magnon is only possible for high-energy magnons above $\Delta_{0}$, such that one decay product can carry opposite spin (cyan). Thus, the lower magnon branch is stable below the gap $\Delta_{0}$. (b) Class-(ii) magnets: Antiferromagnets fluctuate and have spin-degenerate magnons. Stability depends on the curvature of the magnon spectrum: negatively curved spectra ($\kappa<0$) lead to stable magnons, while positively curved spectra ($\kappa>0$) result in instability. (c) Class-(iii) magnets: Altermagnets are characterized by unconventional time-reversal symmetry breaking and beyond-s-wave even-parity spin splitting such as d/g/i-wave. In nodal planes, i.e., directions without spin splitting, the stability mirrors that of class-(ii) magnets, where the curvature of the spectrum dictates whether magnons can decay, but is additionally modified by the altermagnetic splitting, as indicated by the asterisk. Along directions with finite spin splitting the upper magnon branch is universally unstable due to the indicated decay process. The lower magnon branch exhibits rich stability physics that depends on the symmetry of the splitting as indicated by the asterisk: In d-wave magnets, it is either completely stable for $\kappa<0$ or exhibits a d-wave-shaped island of stability in the vicinity of the origin for $\kappa>0$. In contrast, there is no such island of stability for general g-wave and i-wave altermagnets.
@@ Results
For insulators with localized electronic degrees of freedom, all three classes of magnets adhere to the nonrelativistic SO(3)(3)-symmetric Heisenberg model ($\hbar=1$),
@@
where ${\boldsymbol{S}}_{\boldsymbol{r}}$ is a local spin operator of length $S$ at site $\boldsymbol{r}$}, and $J_{{\boldsymbol{r}},{\boldsymbol{r}}^{\prime}}$ the exchange constant between spins at $\boldsymbol{r}$} and $\boldsymbol{r}$$^{\prime}$. In the collinear ground state, spontaneous symmetry breaking reduces SO(3)(3) to SO(2)(2), with the ordered moment chosen along the $z$-axis. Goldstone modes emerge, and the total spin component $S^{z}=\sum_{\boldsymbol{r}}S_{\boldsymbol{r}}^{z}$ remains a good quantum number. Single spin-flip excitations (magnons) thus carry spin $\pm 1$ (also denoted by $\uparrow$ and $\downarrow$) relative to the ground state, and follow dispersions $\epsilon_{\pm}({\boldsymbol{k}})$, where $\boldsymbol{k}$} is the crystal momentum, and $\pm$” indicates spin, not necessarily energetic order. Optical magnon branches are neglected as they do not impact low-energy physics.
@@
Class-(i) magnets: Ferromagnets inherently support stable magnons because their ground state is an exact eigenstate. Ferrimagnets, however, fluctuate because their sublattice-polarized ground state is not an eigenstate. Having two antiparallel but uncompensated sublattice magnetizations, they feature two magnon branches: a gapless branch, $\epsilon_{-}({\boldsymbol{k}})=A|{\boldsymbol{k}}|^{2}$, and a gapped branch, $\epsilon_{+}({\boldsymbol{k}})=A|{\boldsymbol{k}}|^{2}+\Delta_{0}$, where $A>0$ is the spin-wave stiffness and $\Delta_{0}>0$ the gap, see Fig. 1a. Here, s-wave spin splitting manifests as an isotropic energy difference: $\epsilon_{+}({\boldsymbol{k}})-\epsilon_{-}({\boldsymbol{k}})=\Delta_{0}$. The positive curvature of the gapless branch implies that any initial magnon could theoretically decay into three magnons at smaller momenta while conserving energy and momentum. However, Eq. (2) requires at least one decay product to have opposite spin, precluding decay into the gapless branch alone. Decay becomes possible only at energies exceeding $\Delta_{0}$, where one decay product can transition into the gapped branch. Consequently, the gapless magnon branch in class-(i) magnets is universally stable below the energy gap $\Delta_{0}$ and protected by s-wave spin splitting.
@@
Figure 3: Nonlinear spin-wave analysis of a two-dimensional d-wave altermagnet. (a) Schematic of the spin model for a two-dimensional d-wave altermagnet on a checkerboard lattice, described by the Hamiltonian in Eq. (8). The system hosts a collinear $(\pi,\pi)$ Néel order with nearest-neighbor (white bonds) and next-nearest-neighbor (red and blue bonds) exchange interactions $J_{1}$ and $J_{2}\pm\Delta$, respectively. The checkerboard modulation—parametrized by $\Delta$—arises from nonmagnetic atoms (silver spheres) that alter the diagonal exchange, inducing a $d$-wave spin splitting of magnons. (b) Diagrammatic representation of the three-magnon scattering process, the leading mechanism for magnon decay. Due to momentum conservation (up to a reciprocal lattice vector), the sum of the momenta of the three intermediate states equals the initial momentum $\boldsymbol{k}$}. Total spin conservation ensures that an initial single spin-$\uparrow$ magnon scatters into two spin-$\uparrow$ magnons and one spin-$\downarrow$ magnon, with a similar process for an initial spin-$\downarrow$ magnon. (c-f) Renormalized magnon spectrum (lines with dot markers) compared to the linear magnon spectrum (solid lines) for $J_{1}=1$, $J_{2}=-0.5$, $S=1/2$, and selected values of $\Delta$ as indicated. The altermagnetic splitting is most pronounced at $(\frac{\pi}{2},\pm\frac{\pi}{2})$ and zero along the nodal planes, e.g., in the direction from $(0,0)$ to $(\pi,0)$. (g-j) Magnon damping $\Gamma_{\pm}({{\boldsymbol{k}}})$, highlighting the instability of the upper magnon branch. The inset in panel (g) shows the magnetic Brillouin zone with indicated high-symmetry paths. The color of these paths is used on the horizontal momentum axes in panels (c-j) to indicate direction. The cyan/magenta checkerboard pattern indicates the sign of the d-wave altermagnetic spin splitting. The legend in panel (j) distinguishes between nonlinear spin-wave (NLSW) and linear spin-wave (LSW) calculations.
Class-(iii) magnets: We derive general insights into magnon decay in altermagnets by performing a low-energy expansion of the magnon dispersion in two dimensions. We consider an ideal long-wavelength altermagnetic magnon dispersion relation, in which crystallographic symmetries enter only via the altermagnetic splitting, given by
@@
where $v>0$ is the magnon velocity, $\kappa$ represents the leading (cubic) non-altermagnetic nonlinearity, $\tan\varphi_{{\boldsymbol{k}}}=k_{y}/k_{x}$, and $\Delta\geq 0$ parametrizes the altermagnetic splitting with d-wave ($n=2$), g-wave ($n=4$) or i-wave character ($n=6$). The unconventional time-reversal symmetry breaking results in a gap $\epsilon_{+}({\boldsymbol{k}})-\epsilon_{-}({\boldsymbol{k}})=2\Delta\sin\left(n\varphi_{{\boldsymbol{k}}}\right)|{\boldsymbol{k}}|^{n}$. Along a nodal plane (or line in 2D), e.g., ${\boldsymbol{k}}=(k_{x},k_{y})=(k,0)$, the splitting vanishes, suggesting that the class-(ii) curvature argument is relevant. Away from the nodal planes, the magnon bands split, creating novel decay channels, see Fig. 1c.
We first discuss magnon stability along nodal planes. For $\kappa>0$, nodal-plane magnons decay as in class-(ii) magnets. For $\kappa<0$, magnons in class-(ii) magnets remain stable. However, consider the decay depicted in Fig. 2a: a nodal-plane magnon decays into a long-wavelength magnon and two nearby lower-branch magnons. As shown in Supplementary Material I, this process obeys Eq. (2), if $\Delta>\Delta^{\star}$, where
$\Delta^{\star}=\sqrt{3}2^{n-5/2}n^{-1}|{\boldsymbol{k}}|^{2-n}\sqrt{-\kappa}\sqrt{4v+3|{\boldsymbol{k}}|^{2}\kappa}\quad\text{for}\quad\kappa<0.$ (4)
As $|{\boldsymbol{k}}|\to 0$, $\Delta^{\star}$ diverges for $n=4$ (g-wave) and $n=6$ (i-wave), indicating that nodal-plane magnons are stable. For $n=2$ (d-wave), however, $\Delta^{\star}$ converges to
@@
Therefore, d-wave magnets with $\kappa<0$ have unstable nodal-plane magnons if $\Delta>\Delta^{\star}_{\text{d-wave}}$.
Next, we analyze the decay of the lower-branch magnon along a general direction, where it can decay into two lower-branch magnons and a spin-opposite magnon near the Goldstone mode (Fig. 2b). For $\kappa<0$, this decay process is kinematically forbidden. For $\kappa>0$, it remains forbidden if
@@
Thus, in d-wave altermagnets ($n=2$), lower-branch magnons have an “island of stability” at $|{\boldsymbol{k}}|<k^{\star}_{\text{d-wave}}\left|\sin\left(2\varphi_{{\boldsymbol{k}}}\right)\right|$, with
$k^{\star}_{\text{d-wave}}=\frac{2\Delta}{3\kappa}\quad\text{for}\quad\kappa>0.$ (7)
No such solution exists for g-wave magnets ($n=4$) and i-wave magnets ($n=6$).
The upper magnon branch has two decay channels. In Fig. 2c, it decays into a mode close to the origin and two modes to its left and right. This process is possible in d-wave magnets even for $\kappa<0$, as the altermagnetic splitting dominates over the cubic nonlinearity at long wavelengths. The process in Fig. 2d, where it decays into a lower-branch magnon at approximately the same momentum and two upper-branch magnons around the origin, is always kinematically possible. Thus, the upper magnon branch in altermagnets is universally unstable.
@@
To expand on the previous discussion of magnon decay, we consider a d-wave altermagnet on the checkerboard lattice (see Fig. 3a), performing a nonlinear spin-wave analysis. The spin Hamiltonian is
@@
Here, $J_{1}>0$ is the antiferromagnetic nearest-neighbor, and $J_{2}$ the second-nearest-neighbor exchange interaction. $\Delta>0$ represents the d-wave altermagnetic magnon band splitting.
Using a $1/S$ spin-wave expansion around the Néel-ordered ground state based on the Holstein-Primakoff transformation [27], we obtain the noninteracting single-magnon spectrum $\epsilon_{\pm}({\boldsymbol{k}})$ given in Appendix B. At low energies, the spectrum maps onto Eq. (3) with a rotated nodal plane and additional directional dependence of cubic nonlinearities. Up to order $1/S^{2}$, the renormalized spectrum is $\tilde{\epsilon}_{\pm}({\boldsymbol{k}})=\epsilon_{\pm}({\boldsymbol{k}})+\delta\epsilon_{\pm}({\boldsymbol{k}})-\mathrm{i}\Gamma_{\pm}({\boldsymbol{k}})$, where $\delta\epsilon_{\pm}({\boldsymbol{k}})$ corrects the energy and
@@
$\times\delta\left(\epsilon_{\pm}({\boldsymbol{k}})-\epsilon_{\pm}({\boldsymbol{p}})-\epsilon_{\pm}({\boldsymbol{q}})-\epsilon_{\mp}({\boldsymbol{k}}-{\boldsymbol{p}}-{\boldsymbol{q}})\right),$ (9)
@@
is the decay rate associated with the self-energy diagram in Fig. 3b. In Eq. (9), $N_{\textrm{muc}}$ is the number of magnetic unit cells and $W_{\pm}({\boldsymbol{k}},{\boldsymbol{p}},{\boldsymbol{q}})$ is the interaction vertex (Supplementary Material II).
@@
The interacting magnon spectrum $\tilde{\epsilon}_{\pm}({\boldsymbol{k}})$ is shown in Figs. 3c-j for $\kappa<0$, exhibiting d-wave splitting that increases with $\Delta$. Many-body interactions stiffen both bands and cause damping of the upper magnon branch, see Figs. 3g-j. The damping grows with increasing $\Delta$, especially in the short-wavelength limit. When $\Delta>\Delta^{\star}_{\text{d-wave}}$ the damping is nonzero along the nodal planes (Figs. 3h), in agreement with kinematic predictions.
The damping rate $\Gamma_{\pm}({\boldsymbol{k}})$ is related to the density of states $\mathcal{D}_{\pm}(\omega,{\boldsymbol{k}})$ of the three-magnon continuum with net spin $\pm 1$: